Can someone help me solve the solutin of the system using the elemination method?
5x + 2y = -11
7x - 3y = 13

Answer by mananth(16946)

5x+2y=-11.............1
7x-3y=13 .............2

multiply (1)by -3
Multiply (2) by 2
-15x-6y= 33
14x-6y=26
Add the two equations
-x=59
plug value of x in (1)
5x+2y=-11
-295+2y= -11
2y=-11+295
2y=284
y=142